Norway was recently announced as the country with the highest average life expectancy. The average life expectancy is one and half times greater than South Africa, while Germany has an average life expectancy that is 2 years less than the average life expectancy of Norway, South Africa and Germany is 214 years. Determine the life expectancy in each country
1
Expert's answer
2019-05-21T09:33:29-0400

Let average life expectancy of Africa be x, then average life expectancy of Norway is 3/2*x and average life expectancy of Germany is 3/2*x-2, their total life expectancy is x+3/2*x+3/2*x-2.

x+3/2*x+3/2*x-2 = 214

4*x = 216, x = 54

Average life expectancy of Africa is 54 years,

Average life expectancy of Norway is 54*3/2 = 81 years,

Average life expectancy of Germany is 81-2 = 79 years.

Answer: Life expectancy of Africa, Norway and Germany are 54 years, 81 years and 79 years respectively.